# Modeling $(\infty,1)$-categories with Segal spaces

### v1

@@ 3. Construction of the model structure
By the small object argument, the pair (-$(\mathcal{I}\text{-}\mathrm{cof},\mathcal{I}\text{-}\mathrm{fib})$ forms a weak factorization system on $\mathcal{C}$.
@@
[…]
Since (-$(\mathcal{I}\text{-}\mathrm{cof},\mathcal{I}\text{-}\mathrm{fib})$ is a weak factorization system, such $\mathcal{I}$-fibrant replacements always exist.
We are now ready to recall the statement of [GMSV23, Theorem 2.8].
@@
In the remainder of this section, we prove Theorem 3.1 using Theorem 3.4. First, in Sections 3.1 and 3.2, we describe the generating sets $\mathcal{I}$ and $\mathcal{J}$, respectively, and study the weak factorization systems they generate. In particular, we show that $\mathcal{I}$-cofibrations and $\mathcal{J}$-fibrations between $\mathcal{J}$-fibrant objects align with the descriptions provided in (i), (ii), and (iv) of Theorem 3.1. Next, in Section 3.3, we define $\mathcal{W}_{f}$ as the class of Dwyer–Kan equivalences between Segal spaces, obtaining the description given in (iii) of Theorem 3.1. We also verify that, with this definition, Conditions (I) and (II) of Theorem 3.4 are satisfied. Finally, in Sections 3.4, 3.5 and 3.6, we establish that Conditions (III), (IV), and (V) of Theorem 3.4 hold in our setting, thereby completing the proof of Theorem 3.1.
@@ 3.1. Cofibrations and trivial fibrations
In addition, we show that the $\mathcal{I}$-cofibrations coincide with the cofibrations described in (i) of Theorem 3.1.
@@
First note that $\mathcal{A}$ contains all the monomorphisms in $\mathcal{I}$. We show that $\mathcal{A}$ is saturated. The fact that $\mathcal{A}$ is closed under pushouts follows from the fact that $\mathrm{sSet}$ is left proper. To show that $\mathcal{A}$ is closed under transfinite compositions and retracts, we use the characterization (ii) from Lemma 3.7. If $f$ arises as a transfinite composition of maps in $\mathcal{A}$, then each $f_{0}^{-1}(S)\to S$ is either a transfinite composition of trivial cofibrations, hence a weak equivalence, or $f_{0}^{-1}(S)=$ and $S$ is a filtered colimit of weakly contractible complexes, hence weakly contractible. It follows that $f\in\mathcal{A}$ as well. Finally, suppose that we have a retract diagram in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}$ as below left, where $f\in\mathcal{A}$. Given a connected component $S\subseteq W_{0}$, we get an induced retract diagram in $\mathrm{sSet}$ as below right.
$Z$$W$$X$$Y$$Z$$W$$r$$g$$f$$g$$g_{0}^{-1}(S)$$S$$f_{0}^{-1}(r_{0}^{-1}(S))$$r_{0}^{-1}(S)$$g_{0}^{-1}(S)$$S$$g_{0}$$f_{0}$$g_{0}$
@@
We have that $r_{0}^{-1}(S)\subseteq Y_{0}$ is a connected component as the preimage of the connected component $S\subseteq W_{0}$. Hence, by assumption, we have either that $f_{0}^{-1}(r_{0}^{-1}(S))\to r_{0}^{-1}(S)$ is a weak equivalence in $\mathrm{sSet}$ or that $f_{0}^{-1}(r_{0}^{-1}(S))=\emptyset$ and $r_{0}^{-1}(S)\simeq\Delta[0]$. The result then follows from the fact that both of these conditions are closed under retracts. As the class of $\mathcal{I}$-cofibrations is the smallest saturated class containing $\mathcal{I}$, we get that -$\mathcal{I}\text{-}\mathrm{cof}\subseteq\mathcal{A}$.
@@
Conversely, let $f\colon X\hookrightarrow Y$ be a monomorphism in $\mathcal{A}$. Using the small object argument, we factor the map $f$ into an $\mathcal{I}$-cofibration $i\colon X\hookrightarrow Z$ followed by an $\mathcal{I}$-fibration $p\colon Z\to Y$. Note that $f$ and the $\mathcal{I}$-cofibration $i$ are both in $\mathcal{A}$. By Proposition 3.6, the $\mathcal{I}$-fibration $p$ is such that the map $p_{0}\colon Z_{0}\to Y_{0}$ is a fibration in $\mathrm{sSet}$ and the map $p_{0,0}\colon Z_{0,0}\to Y_{0,0}$ is surjective. We claim that $p_{0}$ is in fact a trivial fibration in $\mathrm{sSet}$. It will then follow from Proposition 3.6 that $p$ is a Reedy trivial fibration. Using this and the fact that $f$ is a monomorphism, we get that $f$ is a retract of $i$ and hence an $\mathcal{I}$-cofibration, as desired.
To verify the claim that $p_{0}$ is a trivial fibration in $\mathrm{sSet}$, it remains to show that it is a weak equivalence in $\mathrm{sSet}$. To see this, we prove that it induces a weak equivalence over each path component of the target. Let $S\subseteq Y_{0}$ be a path component and consider the maps
$f_{0}^{-1}(S)=i_{0}^{-1}(p^{-1}_{0}(S))\to p_{0}^{-1}(S)\to S.$
If $f_{0}^{-1}(S)\neq\emptyset$, then both the left-hand map and the composite are weak equivalences in $\mathrm{sSet}$, so that $p_{0}^{-1}(S)\to S$ is also a weak equivalence in $\mathrm{sSet}$. If $f_{0}^{-1}(S)=\emptyset$, then both $S$ and $p_{0}^{-1}(S)$ are weakly contractible, so that $p_{0}^{-1}(S)\to S$ is a weak equivalence in $\mathrm{sSet}$, as well. ∎
@@ 3.2. Anodyne extensions and fibrations between fibrant objects
We now aim to show that $\mathcal{J}$-fibrant objects and $\mathcal{J}$-fibrations between them align with the descriptions given in (ii) and (iv) of Theorem 3.1. For this, we first prove that $\mathcal{J}$-fibrant replacement can be computed as $\mathcal{J}_{0}$-fibrant replacements.
@@
Since the map $F[0]\hookrightarrow NI[1]$ has a retract, every simplicial space has the extension property against $F[0]\hookrightarrow NI[1]$. This shows (i). For (ii), we can apply the small object argument with respect to $\mathcal{J}_{0}$ to provide the desired $\mathcal{J}$-fibrant replacement. ∎
@@
As a $\mathcal{J}$-fibration is by definition a $\mathcal{J}_{0}$-fibration with the right lifting property against $F[0]\hookrightarrow NI[1]$, the result follows from Proposition 2.15, Lemma 3.11 (i), and Remark 3.10. ∎
We now prove a useful property of $\mathcal{J}_{0}$-cofibrations, which can be applied in particular to the $\mathcal{J}$-fibrant replacements from Lemma 3.11.
@@ 3.3. Weak equivalences
Finally, we introduce the class of weak equivalences. To get the desired description of the weak equivalences between fibrant objects stated in (iii) of Theorem 3.1, we define $\mathcal{W}_{f}$ to be the class of Dwyer–Kan equivalences between Segal spaces.
With this definition, we now show that Conditions (I) and (II) of Theorem 3.4.
@@
The class $\mathcal{W}_{f}$ of Dwyer–Kan equivalences between Segal spaces satisfies the $2$-out-of-$6$ property.
@@
Consider the functor $\Phi\colon(\mathrm{sSet}^{\Delta^{\mathrm{op}}})^{[1]}\to\mathrm{Cat}^{[1]}\times\mathrm{sSet}^{[1]}$ sending a map of simplicial spaces $X\to Y$ to the tuple of $cRX\to cRY$ and $X_{1}\to Y_{1}\times_{Y_{0}^{\times 2}}X_{0}^{\times 2}$. We define $\overline{\mathcal{W}}$ to be the inverse image under $\Phi$ of the full subcategory $\mathcal{W}_{\mathrm{Cat}}\times\mathcal{W}_{\mathrm{sSet}}$ of tuples of an equivalence of categories and a weak equivalence in $\mathrm{sSet}$. Both of these are accessible subcategories since they form the weak equivalences of a combinatorial model structure. Since $\Phi$ preserves filtered colimits, it follows that $\overline{\mathcal{W}}$ is an accessible subcategory as well by [Lur09a, Corollary A.2.6.5]. Finally, Lemmas 2.10 and 2.11 imply that a map between Segal spaces $f\colon X\to Y$ is contained in $\overline{\mathcal{W}}$ if and only if it is a Dwyer–Kan equivalence. ∎
@@ 3.4. Path objects
We now show that Condition (III) of Theorem 3.4 holds. This follows from the following result, using Proposition 3.12.
@@
The map $X^{NI[1]}\to X\times X$ is an isofibration by Lemma 2.18, taking $Y=\Delta[0]$. The map $X\to X^{NI[1]}$ is a Dwyer–Kan equivalence by [Rez01, Lemma 13.9]. ∎
@@ 3.5. Isofibrations that are Dwyer–Kan equivalences are trivial fibrations
We now show that Condition (IV) of Theorem 3.4. In light of Proposition 3.12, this condition is precisely the implication “(ii) implies (i)” in the following statement.
@@
In all three cases, the map $f$ is in particular a Reedy fibration. The equivalence between (i) and (iii) then follows from Lemma 2.11 and Proposition 3.6. If $f$ satisfies condition (iii), then $\mathrm{ho}X\to\mathrm{ho}Y$ is fully faithful and surjective on objects, so in particular an isofibration. Hence $f$ is an isofibration and a Dwyer–Kan equivalence. Conversely, suppose that $f$ is an isofibration and a Dwyer–Kan equivalence. To see that $f$ satisfies (iii), the only nontrivial condition to check is that $X_{0,0}\to Y_{0,0}$ is surjective. This follows from the fact that the induced functor $\mathrm{ho}f$ is both an isofibration and an equivalence of categories, so it is surjective on objects. ∎
@@ 3.6. Trivial fibrations are weak equivalences
Finally, we show that Condition (V) of Theorem 3.4 holds. For this, recall the coskeleton functor from Notation 2.8.
@@
Let $p\colon K\to L$ be a fibration in $\mathrm{sSet}$ between Kan complexes such that the induced map $p_{0}\colon K_{0}\to L_{0}$ is surjective, and let $\mathrm{cosk}_{0}(p)\colon\mathrm{cosk}_{0}(K)\to\mathrm{cosk}_{0}(L)$ be the induced map. Then the induced functor
$\mathrm{cosk}_{0}(p)^{*}\colon\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Seg}\ /\mathrm{cosk}_{0}(L)}\to\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Seg}\ /\mathrm{cosk}_{0}(K)}$
is left Quillen.
@@
is left Quillen. To deduce that $\mathrm{cosk}_{0}(p)^{*}\colon\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Seg}\ /\mathrm{cosk}_{0}(L)}\to\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Seg}\ /\mathrm{cosk}_{0}(K)}$ is left Quillen, it suffices to show that it sends the maps $\mathrm{Sp}[m]\hookrightarrow F[m]\to Y$ for $m\geq 2$ to weak equivalences in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Seg}\ /\mathrm{cosk}_{0}(K)}$, by [Hir03, Theorem 3.3.20]. Given $m\geq 0$, consider pullback squares in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}$
@@
To see that $i\colon Q\hookrightarrow P$ is a trivial cofibration in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Seg}}$, it suffices to verify that for any Segal space $X$, the map $i^{*}\colon\mathrm{Map}(P,X)\to\mathrm{Map}(Q,X)$ is a trivial fibration in $\mathrm{sSet}$. To see this, let us write $\alpha_{0},\dots,\alpha_{m}\in L_{0}$ for the images of the vertices of $F[m]$ in $\mathrm{cosk}_{0}(L)_{0,0}\cong L_{0}$. The simplicial space $R$ can then be identified with the constant simplicial simplicial space with value $K_{\alpha_{0}}\amalg\dots\amalg K_{\alpha_{m}}$, where $K_{\alpha_{i}}$ denotes the fiber of $p\colon K\to L$ over $\alpha_{i}\in L_{0}$. Using this and the fact that $X$ is Reedy fibrant, we then obtain a sequence of fibrations in $\mathrm{sSet}$
$\mathrm{Map}(P,X)$$\mathrm{Map}(Q,X)$$\mathrm{Map}(R,X)\cong\mathrm{Map}(K_{\alpha_{0}},X_{0})\times\dots\times\mathrm{Map}(K_{\alpha_{m}},X_{0}).$$i^{*}$
To see that the fibration $i^{*}$ is a trivial fibration, it suffices to verify that it induces a trivial fibration between the fibers over a vertex in $\mathrm{Map}(R,X)$. Let us therefore fix maps of spaces $f_{0}\colon K_{\alpha_{0}}\to X_{0}$, $\dots$, $f_{m}\colon K_{\alpha_{m}}\to X_{0}$. Unraveling the definitions, one sees that the induced map on fibers fits into the following diagram of pullback squares.
@@
Since $X$ is a Segal space, the map $X_{m}\to\mathrm{Map}(\mathrm{Sp}[m],X)$ is a trivial fibration of simplicial sets, hence the map $q$ is a trivial fibration as well. We conclude that $i^{*}$ induces a trivial fibration between the fibers over $(f_{0},\dots,f_{m})$, as desired. ∎
@@
This is straightforward from unpacking the $m$-th relative matching maps and Segal maps, and using Proposition 3.6. ∎
We are now ready to show that Condition (V) of Theorem 3.4 holds.
@@
By Lemma 3.11, we can choose the $\mathcal{J}$-fibrant replacements to be $\mathcal{J}_{0}$-fibrant replacements. Let $j_{Y}\colon Y\to\widetilde{Y}$ be a $\mathcal{J}_{0}$-fibrant replacement and factor the composite $j_{Y}\circ f$ as a $\mathcal{J}_{0}$-cofibration $j_{X}$ followed by a $\mathcal{J}_{0}$-fibration $\widetilde{f}$, as below left. Then, consider the below right commutative diagram in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}$.

### v2

@@ 3. Construction of the model structure
By the small object argument, the pair $(\mathcal{I}\text{-}\mathrm{cof},\mathcal{I}\text{-}\mathrm{fib})$ forms a weak factorization system on $\mathcal{C}$.
@@
    $X~\widetilde{X}Y~\widetilde{Y}$-$\mathcal{I}\text{-}\mathrm{cof}\ni$-$\in\mathcal{I}\text{-}\mathrm{cof}$
Since $(\mathcal{I}\text{-}\mathrm{cof},\mathcal{I}\text{-}\mathrm{fib})$ is a weak factorization system, such $\mathcal{I}$-fibrant replacements always exist.
The following theorem is a combination of [GMSV23, Theorem 2.8 and Proposition 2.21].
@@
In the remainder of this section, we prove Theorem˜3.1 using Theorem˜3.4. First, in Sections˜3.1 and 3.2, we describe the generating sets $\mathcal{I}$ and $\mathcal{J}$, respectively, and study the weak factorization systems they generate. In particular, we show that $\mathcal{I}$-cofibrations and $\mathcal{J}$-fibrations between $\mathcal{J}$-fibrant objects align with the descriptions provided in (i), (ii), and (iv) of Theorem˜3.1. Next, in Section˜3.3, we define $\mathcal{W}_{f}$ as the class of Dwyer–Kan equivalences between Segal spaces, obtaining the description given in (iii) of Theorem˜3.1. We also verify that, with this definition, Conditions (I) and (II) of Theorem˜3.4 are satisfied. Finally, in Sections˜3.4, 3.5 and 3.6, we establish that Conditions (III), (IV), and (V) of Theorem˜3.4 hold in our setting, thereby completing the proof of Theorem˜3.1.
@@ 3.1. Cofibrations and trivial fibrations
In addition, we show that the $\mathcal{I}$-cofibrations coincide with the cofibrations described in (i) of Theorem˜3.1.
@@
First note that $\mathcal{A}$ contains all the monomorphisms in $\mathcal{I}$. We show that $\mathcal{A}$ is saturated. The fact that $\mathcal{A}$ is closed under pushouts follows from the fact that taking pushout along a map in $\mathrm{sSet}$ induces a left Quillen functor between slices. To show that $\mathcal{A}$ is closed under transfinite compositions and retracts, we use the characterization (ii) from Lemma˜3.7. If $f$ arises as a transfinite composition of maps in $\mathcal{A}$, then each $f_{0}^{-1}(S)\to S$ is either a transfinite composition of trivial cofibrations, hence a weak equivalence, or $f_{0}^{-1}(S)=\emptyset$ and $S$ is a filtered colimit of weakly contractible complexes, hence weakly contractible. It follows that $f\in\mathcal{A}$ as well. Finally, suppose that we have a retract diagram in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}$ as below left, where $f\in\mathcal{A}$. Given a connected component $S\subseteq W_{0}$, we get an induced retract diagram in $\mathrm{sSet}$ as below right.
[…]
@@
Here we write $T\subseteq r_{0}^{-1}(S)\subseteq Y_{0}$ for the connected component of $Y_{0}$ containing the image $i_{0}(S)$ of the connected component $S\subseteq W_{0}$. By assumption, we have either that $f_{0}^{-1}(T)\to T$ is a weak equivalence in $\mathrm{sSet}$ or that $f_{0}^{-1}(T)=\emptyset$ and $T\simeq\Delta[0]$. The result then follows from the fact that both of these conditions are closed under retracts. As the class of $\mathcal{I}$-cofibrations is the smallest saturated class containing $\mathcal{I}$, we get that -$\mathcal{I}\text{-}\mathrm{cof}\subseteq\mathcal{A}$.
Conversely, let $f\colon X\hookrightarrow Y$ be a monomorphism in $\mathcal{A}$. Let $R$ be the set of path components of $Y_{0}$ that are not in the image of $f_{0}$, and choose a map $R\to Y_{0}$ selecting exactly one vertex in each such path component. The map $f$ then factors as an $\mathcal{I}$-cofibration $X\hookrightarrow X\amalg R$ followed by a monomorphism $f^{\prime}\colon X\amalg R\hookrightarrow Y$ such that $f^{\prime}_{0}$ is a weak equivalence; in particular, $f^{\prime}$ is also in $\mathcal{A}$, and it suffices to verify that it is contained in -$\mathcal{I}\text{-}\mathrm{cof}$.
@@
We now apply the small object argument with respect to the set of maps $\mathcal{I}^{\prime}=\mathcal{I}\setminus\{\emptyset\to F[0]\}$, in order to factor the map $f^{\prime}$ into an $\mathcal{I}^{\prime}$-cofibration $i\colon X\amalg R\hookrightarrow Z$ followed by an $\mathcal{I}^{\prime}$-fibration $p\colon Z\to Y$. We claim that $p$ is a Reedy trivial fibration. Using this and the fact that $f^{\prime}$ is a monomorphism, we get that $f^{\prime}$ is a retract of $i$. It follows that $f^{\prime}$ is an $\mathcal{I}^{\prime}$-cofibration and hence an $\mathcal{I}$-cofibration.
To see that $p$ is a Reedy trivial fibration, note that the right lifting property against the maps in $\mathcal{I}^{\prime}$ implies that $p$ satisfies properties (1) and (2) of Proposition˜3.6. It therefore remains to check that $p_{0}\colon Z_{0}\to Y_{0}$ is a weak equivalence in $\mathrm{sSet}$, or equivalently, by the 2-out-of-3 property, that $i_{0}\colon X_{0}\amalg R\to Z_{0}$ is a weak equivalence. This follows from that fact that every map in $\mathcal{I}^{\prime}$ induces a trivial cofibration in $\mathrm{sSet}$ in degree 0, so that $i_{0}\colon X_{0}\amalg R\to Z_{0}$ is a trivial cofibration in $\mathrm{sSet}$. ∎
@@ 3.2. Anodyne extensions and fibrations between fibrant objects
We now aim to show that $\mathcal{J}$-fibrant objects and $\mathcal{J}$-fibrations between them align with the descriptions given in (ii) and (iv) of Theorem˜3.1. For this, we first prove that $\mathcal{J}$-fibrant replacement can be computed as $\mathcal{J}_{0}$-fibrant replacements.
@@
Since the map $F[0]\hookrightarrow NI[1]$ has a retraction, every simplicial space has the extension property against $F[0]\hookrightarrow NI[1]$. This shows (i). For (ii), we can apply the small object argument with respect to $\mathcal{J}_{0}$ to provide the desired $\mathcal{J}$-fibrant replacement. ∎
@@
As a $\mathcal{J}$-fibration is by definition a $\mathcal{J}_{0}$-fibration with the right lifting property against $F[0]\hookrightarrow NI[1]$, the result follows from Proposition˜2.15, Lemma˜3.11 (i), and Remark˜3.10. ∎
We now prove a useful property of $\mathcal{J}_{0}$-cofibrations, which can be applied in particular to the $\mathcal{J}$-fibrant replacements from Lemma˜3.11.
@@ 3.3. Weak equivalences
Finally, we introduce the class of weak equivalences. To get the desired description of the weak equivalences between fibrant objects stated in (iii) of Theorem˜3.1, we define $\mathcal{W}_{f}$ to be the class of Dwyer–Kan equivalences between Segal spaces.
With this definition, we now show that Conditions (I) and (II) of Theorem˜3.4 hold.
@@
The class $\mathcal{W}_{f}$ of Dwyer–Kan equivalences between Segal spaces satisfies the 22-out-of-66 property.
@@
Consider the functor $\Phi\colon(\mathrm{sSet}^{\Delta^{\mathrm{op}}})^{[1]}\to\mathrm{Cat}^{[1]}\times\mathrm{sSet}^{[1]}$ sending a map of simplicial spaces $X\to Y$ to the tuple of $cRX\to cRY$ and $X_{1}\to Y_{1}\times_{Y_{0}^{\times 2}}X_{0}^{\times 2}$. We define $\overline{\mathcal{W}}$ to be the inverse image under $\Phi$ of the full subcategory $\mathcal{W}_{\mathrm{Cat}}\times\mathcal{W}_{\mathrm{sSet}}$ of tuples of an equivalence of categories and a weak equivalence in $\mathrm{sSet}$. Both of these are accessible subcategories since they form the weak equivalences of a combinatorial model structure. Since $\Phi$ preserves filtered colimits, it follows that $\overline{\mathcal{W}}$ is an accessible subcategory as well by [Lur09a, Corollary A.2.6.5]. Finally, Lemmas˜2.10 and 2.11 imply that a map between Segal spaces $f\colon X\to Y$ is contained in $\overline{\mathcal{W}}$ if and only if it is a Dwyer–Kan equivalence. ∎
@@ 3.4. Path objects
We now show that Condition (III) of Theorem˜3.4 holds. This follows from the following result, using Proposition˜3.12.
@@
The map $X^{NI[1]}\to X\times X$ is an isofibration by Lemma˜2.18, taking $Y=\Delta[0]$. The map $X\to X^{NI[1]}$ is a Dwyer–Kan equivalence by [Rez01, Lemma 13.9]. ∎
@@ 3.5. Isofibrations that are Dwyer–Kan equivalences are trivial fibrations
We now show that Condition (IV) of Theorem˜3.4 holds. In light of Proposition˜3.12, this condition is precisely the implication “(ii) implies (i)” in the following statement.
@@
In all three cases, the map $f$ is in particular a Reedy fibration. The equivalence between (i) and (iii) then follows from Lemma˜2.11 and Proposition˜3.6. If $f$ satisfies condition (iii), then $\mathrm{ho}X\to\mathrm{ho}Y$ is fully faithful and surjective on objects, so in particular an isofibration. Hence $f$ is an isofibration and a Dwyer–Kan equivalence. Conversely, suppose that $f$ is an isofibration and a Dwyer–Kan equivalence. To see that $f$ satisfies (iii), the only nontrivial condition to check is that $X_{0,0}\to Y_{0,0}$ is surjective. This follows from the fact that the induced functor $\mathrm{ho}f$ is both an isofibration and an equivalence of categories, so it is surjective on objects. ∎
@@ 3.6. Trivial fibrations are weak equivalences
Finally, we show that Condition (V) of Theorem˜3.4 holds. For this, recall the coskeleton functor from Notation˜2.8.
@@
Let $p\colon K\to L$ be a fibration in $\mathrm{sSet}$ between Kan complexes such that the induced map $p_{0}\colon K_{0}\to L_{0}$ is surjective. Then the functor
$\mathrm{cosk}_{0}(p)^{*}\colon\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Seg}\ /\mathrm{cosk}_{0}(L)}\to\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Seg}\ /\mathrm{cosk}_{0}(K)},$
obtained by taking pullback along the map $\mathrm{cosk}_{0}(p)\colon\mathrm{cosk}_{0}(K)\to\mathrm{cosk}_{0}(L)$, is left Quillen.
@@
is left Quillen. To deduce that $\mathrm{cosk}_{0}(p)^{*}\colon\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Seg}\ /\mathrm{cosk}_{0}(L)}\to\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Seg}\ /\mathrm{cosk}_{0}(K)}$ is left Quillen, it suffices to show that it sends the inner horn inclusions $L^{k}[m]\hookrightarrow F[m]\to\mathrm{cosk}_{0}(L)$ for $0<k<m$ to weak equivalences in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Seg}\ /\mathrm{cosk}_{0}(K)}$, by [Hir03, Theorem 3.3.19] and [JT07, Lemma 3.5]. Given $0<k<m$, consider the following pullback squares in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}$
@@
Let us first show that $i$ is a trivial cofibration in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Seg}}$ when $K$ and $L$ are sets, so that all objects in the above diagram are contained in the subcategory $\mathrm{Set}^{\Delta^{\mathrm{op}}}\subseteq\mathrm{sSet}^{\Delta^{\mathrm{op}}}$. Let us write $K(i)=p^{-1}(\alpha(i))$ for the fibre of $p$ over the image of the $i$-th vertex of $F[m]$ in $\mathrm{cosk}_{0}(L)_{0}=L$. Then $P=N\mathcal{C}$ is the nerve of the category with:
  • objects given by tuples $(i,x)$ with $0\leq i\leq m$ and $x\in K(i)$.
  • exactly one morphism $(i,x)\to(j,y)$ whenever $i\leq j$, and no morphism otherwise.
Let $\sigma\colon(i_{0},x_{0})\to\dots\to(i_{n},x_{n})$ be an $n$-simplex in $P$. Then $\sigma$ is not contained in $Q$ if and only if $\{0,\dots,\hat{k},\dots,m\}\subseteq\{i_{0},\dots,i_{n}\}$. Given such a non-degenerate chain that is not contained in $Q$, let us define its pivot to be the first element $(i_{a},x_{a})$ in the chain with $i_{a}=k$; note that $\sigma$ may not have a pivot.
Let us now fix an element $z\in K(k)$ in the (non-empty) fibre $K(k)$ and say that a non-degenerate chain $\sigma$ not contained in $Q$ is:
  • bounding if it has a pivot of the form $(k,x_{a}=z)$, and
  • bounded if not. In particular, a chain without any term of the form $(k,x)$ is bounded.
Every bounded chain $\sigma$ is an inner face of exactly one bounding chain $\tau$, obtained from $\sigma$ by adding a pivot $(k,z)$ as follows, where $i_{a}=k$ or $i_{a}=k+1$.
@@
Conversely, if $\tau$ is a (by definition non-degenerate) bounding chain with pivot $(i_{a},x_{a})=(k,z)$, then $d_{j}(\tau)$ is a lower-dimensional bounding chain or contained in $P$ when $j\neq a$, while the inner face $d_{a}(\tau)$ is a bounded chain of lower dimension. It follows from this that $Q\hookrightarrow P$ is an iterated pushout of inner horn inclusions: proceeding by induction on the dimension, we can take pushouts along $L^{a}[n]\to F[n]$ to add a bounding chain $\tau$ of length $n$, together with its (unique) bounded inner face $d_{a}(\tau)$. Since every iterated pushout of inner horn inclusions is a trivial cofibration in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Seg}}$ by [JT07, Lemma 3.5], it follows that $Q\hookrightarrow P$ is a trivial cofibration.
Finally, let us treat the case where $p\colon K\to L$ is a map of simplicial sets. Since $p$ is a Kan fibration and surjective on vertices, each map $p_{n}\colon K_{n}\to L_{n}$ is surjective. For $n\geq 0$, we regard $P_{-,n}$ and $Q_{-,n}$ as objects of $\mathrm{sSet}^{\Delta^{\mathrm{op}}}$ through the canonical inclusion $\mathrm{Set}^{\Delta^{\mathrm{op}}}\hookrightarrow\mathrm{sSet}^{\Delta^{\mathrm{op}}}$. Applying the previous argument, the induced map
$Q_{-,n}=L^{k}[m]\times_{\mathrm{cosk}_{0}(L_{n})}\mathrm{cosk}_{0}(K_{n})\to F[m]\times_{\mathrm{cosk}_{0}(L_{n})}\mathrm{cosk}_{0}(K_{n})=P_{-,n}$
is a weak equivalence in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Seg}}$. Moreover, these maps assemble into a natural transformation of functors $\Delta^{\mathrm{op}}\to\mathrm{sSet}^{\Delta^{\mathrm{op}}}$ and the induced map $\mathrm{hocolim}_{[n]\in\Delta^{\mathrm{op}}}Q_{-,n}\to\mathrm{hocolim}_{[n]\in\Delta^{\mathrm{op}}}P_{-,n}$ between homotopy colimits is also a weak equivalence in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Seg}}$. Since $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Seg}}$ is a simplicial model structure for the levelwise tensoring over $\mathrm{sSet}$, these homotopy colimits can be computed by the diagonal by [Hir03, Theorem 15.11.6 and Theorem 18.7.6]. Using that $P_{-,n}$ and $Q_{-,n}$ are constant in the space direction, we conclude that $i\colon Q\hookrightarrow P$ is a weak equivalence, as desired. ∎
@@
  2. (ii)
    Let $f\colon X\to Y$ be an $\mathcal{I}$-fibration. Then the map $X\to Y\times_{\mathrm{cosk}_{0}(Y_{0})}\mathrm{cosk}_{0}(X_{0})$ is a Reedy trivial fibration.
@@
This is straightforward from unpacking the $m$-th relative matching maps and Segal maps, and using Proposition˜3.6. The $m$-th relative matching map of $X\to Y\times_{\mathrm{cosk}_{0}(Y_{0})}\mathrm{cosk}_{0}(X_{0})$ is isomorphic to the $m$-th matching map of $X\to Y$ if $m\geq 1$, and an isomorphism when $m=0$. ∎
We are now ready to show that Condition (V) of Theorem˜3.4 holds.
@@
By Lemma˜3.11, we can choose the $\mathcal{J}$-fibrant replacements to be $\mathcal{J}_{0}$-fibrant replacements. Let $j_{Y}\colon Y\to\widetilde{Y}$ be a $\mathcal{J}_{0}$-fibrant replacement and factor the composite $j_{Y}\circ f$ as a $\mathcal{J}_{0}$-cofibration $j_{X}$ followed by a $\mathcal{J}_{0}$-fibration $\widetilde{f}$, as below left. Then, consider the below right commutative diagram in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}$.